OpenStudy (anonymous):

if f(x)=(x+1)(1-x^2), evaluate f(a+3)

OpenStudy (anonymous):

some help me

OpenStudy (anonymous):

Plug in (a+3) for every value of x in f(x).

OpenStudy (anonymous):

uh huh i got =((a+3)+1)(1-(a+3)^2) dont know what to do after that

OpenStudy (anonymous):

\[f(a+3) = (a+3+1)(1-(a+3)^2)\] just like pjschlotter said.

OpenStudy (anonymous):

Got it!

OpenStudy (anonymous):

Solve it like: (a+3+1) = (a+4) (a+4) (1-(a^2 +6a + 9)) (a+4) (-a^2 -6a -8)

OpenStudy (anonymous):

You can also multiply (a+4) for the rest, but I don't think it's necessary... this form is clearer.

OpenStudy (anonymous):

thank you where is the 6a from?

OpenStudy (anonymous):

It's from the development of (a+3)^2, right? (b+c)^2 = b^2 + 2.b.c + c^2 So, (a+3)^2 = a^2 + 2.a.3 + 3^2 = a^2 + 6a + 3^2

OpenStudy (anonymous):

ok kinda got it so if the number is 6 instead of 3 then it would be 12a instead of 6a correct?

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

so is the solution -12a^2-32a-32

OpenStudy (anonymous):

It's (a+4) (-a^2 -6a -8), as I said!

OpenStudy (anonymous):

oh i though you have to combine them. but thanks alot how do i become a fan?

OpenStudy (anonymous):

can you help me with another problem?

OpenStudy (anonymous):

Of course, just send it. :)

OpenStudy (anonymous):

Erm, actually, I gotta go sleep. lol Class tomorrow. See ya.